OpenStudy (anonymous):

A triangular prism and a triangular pyramid have the same base and height. What is true about the volumes of the triangular prism and the triangular pyramid? The triangular pyramid has a volume one-third the volume of the triangular prism. The triangular pyramid has a volume three times the volume of the triangular prism. The triangular pyramid has a volume one-half the volume of the triangular prism. The triangular prism has a volume one-third the volume of the triangular pyramid.

OpenStudy (danjs):

.

Directrix (directrix):

The volume of the triangular prism is found by the area of its B (base area) times the height of the prism. The volume of the triangular pyramid is found by take 1/3 its B (base area) times the height of the pyramid. The two solids have the same base area and the same height. What can you conclude? @Acecool

Directrix (directrix):

Algebraically, Vprism = B * h Vpyramid = (1/3) * B * h Recall that in this problem, the two solids have the same Base and the same height.